OpenStudy (anonymous):

Jason has two bags with 6 tiles each. The tiles in each bag are shown below: Six squares are numbered sequentially from 1 to 6. Without looking, Jason draws a tile from the first bag and then a tile from the second bag. What is the probability of Jason drawing the tile numbered 5 from the first bag and an odd tile from the second bag? 3 over 6 4 over 6 3 over 36 4 over 36

OpenStudy (phi):

can you answer this part What is the probability of Jason drawing the tile numbered 5 from the first bag ? you want the fraction: # of 5's in the bag divided by the total number of tiles in the bag

OpenStudy (anonymous):

would my answer be 3/36 @phi

OpenStudy (phi):

can you answer What is the probability of Jason drawing the tile numbered 5 from the first bag ?

OpenStudy (anonymous):

but the probability is 5/6

OpenStudy (phi):

the probability of picking the number "5" from a bag with 1,2,3,4,5 in it is not 5/6 the top number in the fraction is the "number of 5's in the bag" how many cards are labeled 5?

OpenStudy (anonymous):

1/5

OpenStudy (phi):

yes. the idea is if you had lots of people pick a card from the bag (then put it back) about 1/5 people would have picked the 5. now we find the probability of picking an odd number from the bag: 1,2,3,4,5 how many of those are odd numbers?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

sry i take so long to reply :b

OpenStudy (phi):

yes, so the probability of picking an odd number is # of odd numbers divided by total number of cards in the bag 3/6

OpenStudy (phi):

the probability of both picking a 5 from the first bag and an odd number from the 2nd bag is the product (i.e. multiply) of the two probabilities (I guess you have to memorize this, as I can't explain it easily) so your answer is 1/6 * 3/6 multiply top time top and bottom times bottom

OpenStudy (phi):

some of this should make a little sense. for example, if you had 1,3,5,7,9 in the bag, the chances of picking an odd number is 100% (they are all odd) and if we calculate # of odd divided by number of cards we get 5/5= 1 (or 100%) on the other hand if we had 2,4,6,8,10 the chances of getting an odd is 0 if we calculate it, # of odds is 0 and 0/5= 0 so at least those cases make sense, right?

OpenStudy (anonymous):

right, so my answer is 3/6

OpenStudy (phi):

your answer is \[ \frac{1}{6} \cdot \frac{3}{6} \] to multiply fractions, multiply top times top and bottom times bottom

OpenStudy (anonymous):

oh ok so my answer is 3/36 right?

OpenStudy (phi):

yes. 3/36 (or 1/12 if they simplify it... sometimes they will do that to trick you)

OpenStudy (anonymous):

ok then, thank you !!! ^_^